# GI-NAS: Boosting Gradient Inversion Attacks through Adaptive Neural Architecture Search

### v3

@@ Abstract
Index Terms:
Federated learning, gradient inversion attacks, neural architecture search, privacy leakage.
@@ I Introduction
Model $1$
Model $2$
Model $3$
Model $4$
Model $5$
@@
Figure 1: Quantitative and qualitative results when randomly attacking $3$ different batches by $5$ different models on ImageNet. In Figure 1(b), the first, second, and third rows are respectively from Batch $1$, Batch $2$, and Batch $3$.
@@
Inspired by the above phenomenons, we propose a novel gradient inversion method, named Gradient Inversion via Neural Architecture Search (GI-NAS), to better match each batch with the optimal model architecture. Specifically, we first enlarge the potential search space for the over-parameterized network by designing different upsampling modules and skip connection patterns. To reduce the computational overhead, we utilize a training-free search strategy that compares the initial gradient matching loss for a given batch over all the candidates and selects the best of them for the final optimization. We further provide substantial experimental evidence that such a metric highly correlates with the real attack performance. We also consider more rigorous and realistic scenarios where the victims may hold high-resolution images and large-sized batches for training, and evaluate advanced defense strategies. Extensive experiments validate that GI-NAS can achieve state-of-the-art performance compared to existing gradient inversion methods. To the best of our knowledge, we are the first to introduce Neural Architecture Search (NAS) to Gradient Inversion Attacks. Our main contributions are as follows:
@@ III Problem Formulation
Basics of Gradient Inversion. We consider the training process of a classification model $f_{\theta}$ parameterized by $\theta$ in FL scenarios. The real gradients $\mathbf{g}$ are calculated from a private batch (with real images $\mathbf{x}$ and real labels $\mathbf{y}$) at the client side. The universal goal of Gradient Inversion Attacks is to search for some fake images $\mathbf{\hat{x}}\in\mathbb{R}^{B\times H\times W\times C}$ with labels $\mathbf{\hat{y}}\in\{0,1\}^{B\times L}$ so that ($\mathbf{\hat{x}}$, $\mathbf{\hat{y}}$) can be close to ($\mathbf{x}$, $\mathbf{y}$) as much as possible, where $B$, $H$, $W$, $C$, and $L$ are respectively batch size, image height, image width, number of channels, and number of classes. This can be realized by minimizing the gradient matching loss [9]:
@@
where $F(\mathbf{\hat{x}})=\nabla_{\theta}\mathcal{L}(f_{\theta}(\mathbf{\hat{x}}),\mathbf{\hat{y}})$ calculates the gradients of $f_{\theta}$ provided with $\mathbf{\hat{x}}$.
@@
$\mathbf{\hat{x}^{*}}=\mathop{\arg\min}_{\mathbf{\hat{x}}}\mathcal{D}(F(\mathbf{\hat{x}}),\mathbf{g})+\lambda\mathcal{R}_{prior}(\mathbf{\hat{x}}),$ (3)
@@ IV-B Stage 1: Training-free Optimal Model Search
  • Input for this stage: Real gradients $\mathbf{g}$, fixed latent code $\mathbf{z_{0}}$ sampled from $\mathcal{N}(0,1)$, and the model search space =$\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$.
@@ IV-B1 Model Search Space Design
Search Space for Upsampling Modules. We decompose the upsampling operations into five key components: feature upsampling, feature transformation, activation function, kernel size, and dilation rate. Then, we allocate a series of possible options to each of these components. When deciding on feature upsampling, we choose from commonly used interpolation techniques, such as bilinear interpolation, bicubic interpolation, and nearest-neighbour interpolation. As for feature transformation, we choose from classical convolution techniques, such as 2D convolution, separable convolution, and depth-wise convolution. As regards activation function, we select from ReLU, LeakyReLU, PReLU, etc. Furthermore, we supply kernel size and dilation rate with more choices, such as $1\times 1$, $3\times 3$, or $5\times 5$ for kernel size and $1$, $3$, or $5$ for dilation rate. The combination of these flexible components can contribute to the diversity of upsampling modules.
Search Space for Skip Connection Patterns. We assume that there are $t$ levels of encoders and decoders in total, and denote them as $e_{1},e_{2},...,e_{t}$ and $d_{1},d_{2},...,d_{t}$. As shown in Figure 3, We consider different skip connection patterns between encoders and decoders. To represent each of these patterns, we define a skip connection matrix $\mathbf{A}\in\{0,1\}^{t\times t}$ that serves as a mask to determine whether there will be new residual connections [68] between pairs of encoders and decoders. More concretely, $\mathbf{A}_{ij}=1$ indicates that there exists a skip connection from $e_{i}$ to $d_{j}$ and $\mathbf{A}_{ij}=0$ means that there is not such a skip connection. As the shapes of feature maps across different network levels can vary significantly (e.g., $64\times 64$ for the output of $e_{i}$ and $256\times 256$ for the input of $d_{j}$), we introduce connection scale factors to tackle this inconsistency and decompose all the possible scale factors into a series of 2×2$\times$2 × upsampling operations or downsampling operations with shared weights. By allocating 00 or $1$ to each of the $t^{2}$ bits in A, we broaden the search space of skip connection patterns as there are $2^{t^{2}}$ possibilities altogether and we only need to sample a portion of them.
@@ IV-B2 Optimal Model Selection
We build up our model search space $\mathcal{M}$ by combining the possibilities of the aforementioned upsampling modules and skip connection patterns. We assume that the size of our model search space is $n$ and the candidates inside it are denoted as =$\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$. We first sample the latent code $\mathbf{z_{0}}$ from the Gaussian distribution and freeze its values on all the models for fair comparison. We then traverse the model search space and calculate the initial gradient matching loss of each individual $G_{r}\,(1\leq r\leq n)$:
@@ IV-C Stage 2: Private Batch Recovery via the Optimal Model
2:$\mathbf{\hat{x}^{*}}$: the generated images;
3:Build up the model search space: =$\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$ with the parameters $\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$
4:$\mathbf{z_{0}}\leftarrow\mathcal{N}(0,1)$, $loss_{min}\leftarrow+\infty$
@@ V-A Experimental Setup
Evaluation Settings. We test our method on CIFAR-10 [69] and ImageNet [22] with the resolution of $32\times 32$ and $256\times 256$. Unlike many previous methods [18, 12] that scale down the ImageNet images to $64\times 64$, here we emphasize that we adopt the high-resolution version of ImageNet. Thus, our settings are more rigorous and realistic. Following previous gradient inversion works [9, 13], we adopt ResNet-$18$ [68] as our global model and utilize the same preprocessing procedures. We build up the search space and randomly generate the alternative models by arbitrarily changing the options of upsampling modules and skip connection patterns.
@@ V-D Further Analysis
TABLE IX: Attack performance when increasing the training iterations for the search metric on ImageNet ($256\times 256$) with the default batch size $B=4$. Note that the training iteration of 00 indicates the use of our training-free search metric.
@@
Ablation Study on Search Metric Training Iteration. GI-NAS utilizes the initial gradient matching loss as the training-free search metric, which implies that its training iteration for each candidate is 00. To explore the rationality behind this design, we gradually increase the training iterations for the search metric. From Table IX, we conclude that further increasing the training iterations cannot improve the attacks and will instead increase the time costs. Hence, utilizing the initial gradient matching loss to find the optimal models is reasonable and computationally efficient.
@@
Dataset Model Random Seed Initial Gradient Matching Loss (×$\times 10^{-3}$) $\sigma_{loss}$ (×$\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{C1}$ $Seed_{C1}$ 8.5647 0.0012 28.7283 0.1211 $Seed_{C2}$ 8.5648 28.8082 $Seed_{C3}$ 8.5645 28.5703 $G_{C2}$ $Seed_{C1}$ 8.4321 0.0006 33.5848 0.2581 $Seed_{C2}$ 8.4323 33.4631 $Seed_{C3}$ 8.4322 33.9584 $G_{C3}$ $Seed_{C1}$ 7.9501 0.0015 49.4290 0.2936 $Seed_{C2}$ 7.9499 49.1171 $Seed_{C3}$ 7.9502 49.7039 ImageNet $G_{I1}$ $Seed_{I1}$ 21.0167 5.6866 22.9013 0.0329 $Seed_{I2}$ 22.1401 22.8912 $Seed_{I3}$ 21.4250 22.8398 $G_{I2}$ $Seed_{I1}$ 24.4600 2.2089 21.1475 0.3040 $Seed_{I2}$ 24.0722 21.6449 $Seed_{I3}$ 24.0828 21.6989 $G_{I3}$ $Seed_{I1}$ 18.5319 6.0290 23.5278 0.1884 $Seed_{I2}$ 18.3819 23.2974 $Seed_{I3}$ 17.4208 23.6709
@@
Dataset Model Random Seed Initial Gradient Matching Loss (×$\times 10^{-3}$) $\sigma_{loss}$ (×$\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{C4}$ $Seed_{C4}$ 14.5338 0.0018 26.4506 0.1925 $Seed_{C5}$ 14.5335 26.2082 $Seed_{C6}$ 14.5336 26.0703 $G_{C5}$ $Seed_{C4}$ 13.4751 0.0018 30.1240 0.2675 $Seed_{C5}$ 13.4753 30.2045 $Seed_{C6}$ 13.4755 30.6222 $G_{C6}$ $Seed_{C4}$ 13.2704 0.0015 31.3139 0.0566 $Seed_{C5}$ 13.2701 31.4256 $Seed_{C6}$ 13.2702 31.3540 ImageNet $G_{I4}$ $Seed_{I4}$ 23.1931 4.6280 21.8376 0.1164 $Seed_{I5}$ 23.6335 22.0556 $Seed_{I6}$ 22.7083 21.8760 $G_{I5}$ $Seed_{I4}$ 19.5264 1.1102 23.6033 0.1889 $Seed_{I5}$ 19.5677 23.2439 $Seed_{I6}$ 19.7360 23.3226 $G_{I6}$ $Seed_{I4}$ 25.9479 3.6054 20.4730 0.1655 $Seed_{I5}$ 25.3497 20.8007 $Seed_{I6}$ 25.9975 20.6776
@@
Preferences of NAS Outcomes. To figure out the implications behind NAS preferences, we trace back the searched optimal models from Section V-C and analyze their averaged architectural statistics in Table XIII. Note that the bicubic operation takes more neighbourhood pixels than the bilinear operation ($16>4$) and is thus considered to be more complex. Besides, the pixel shuffle operation rearranges lots of pixels and is also considered to be more complicated than the bilinear operation. From Table XIII, we observe that the optimal models are largely dependent on recovery difficulties. Generally, batches with larger sizes or under defenses are more difficult to recover. Compared to the default settings (highlighted in gray color) that are with the batch size $B=4$ and under no defenses, the more difficult settings would prefer models with higher complexity (e.g., more skip connections, larger kernel sizes, more complex operations). The complexity gains $\Delta_{connection}$ and $\Delta_{upsampling}$ are often greater than 00 when the settings are $B>4$ or under defenses. Easier batches prefer simpler, lighter architectures and harder batches prefer more complex, heavier architectures. This might be used to further improve the efficiency. By considering the difficulties ahead of time, we can have a better estimation of the complexity of optimal models. Thus, we only need to focus on models with a certain type of complexity and skip other unrelated models even when there is not enough time for NAS.

### v4

@@ I Introduction
Model 11
Model 22
Model 33
Model 44
Model 55
@@
Figure 1: Quantitative and qualitative results when randomly attacking 33 different batches by 55 different models on ImageNet. In Figure 1(b), the first, second, and third rows are respectively from Batch $1$, Batch $2$, and Batch 33.
@@
Inspired by the above phenomena, we propose a novel gradient inversion method, named Gradient Inversion via Neural Architecture Search (GI-NAS), to better match each batch with the optimal model architecture. Specifically, we first enlarge the potential search space for the over-parameterized network by designing different upsampling modules and skip connection patterns. To reduce the computational overhead, we utilize a training-free search strategy that compares the initial gradient matching loss for a given batch over all the candidates and selects the best of them for the final optimization. We further provide substantial experimental evidence that such a metric highly correlates with the real attack performance. We also consider more rigorous and realistic scenarios where the victims may hold high-resolution images and large-sized batches for training, and evaluate advanced defense strategies. Extensive experiments validate that GI-NAS can achieve state-of-the-art performance compared to existing gradient inversion methods. To the best of our knowledge, we are the first to introduce Neural Architecture Search (NAS) to Gradient Inversion Attacks. Our main contributions are as follows:
@@ III Problem Formulation
[…]
@@
[…]
@@
[…]
@@ IV-B Stage 1: Training-free Optimal Model Search
  • Input for this stage: Real gradients $\mathbf{g}$, fixed latent code $\mathbf{z_{0}}$ sampled from $\mathcal{N}(0,1)$, and the model search space $\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$.
@@ IV-B1 Model Search Space Design
Search Space for Upsampling Modules. We decompose the upsampling operations into five key components: feature upsampling, feature transformation, activation function, kernel size, and dilation rate. Then, we allocate a series of possible options to each of these components. When deciding on feature upsampling, we choose from commonly used interpolation techniques, such as bilinear interpolation, bicubic interpolation, and nearest-neighbour interpolation. As for feature transformation, we choose from classical convolution techniques, such as 2D convolution, separable convolution, and depth-wise convolution. As regards activation function, we select from ReLU, LeakyReLU, PReLU, etc. Furthermore, we supply kernel size and dilation rate with more choices, such as $1\times 1$, $3\times 3$, or $5\times 5$ for kernel size and $1$, 33, or $5$ for dilation rate. The combination of these flexible components can contribute to the diversity of upsampling modules.
Search Space for Skip Connection Patterns. We assume that there are $t$ levels of encoders and decoders in total, and denote them as $e_{1},e_{2},...,e_{t}$ and $d_{1},d_{2},...,d_{t}$. As shown in Figure 3, We consider different skip connection patterns between encoders and decoders. To represent each of these patterns, we define a skip connection matrix $\mathbf{A}\in\{0,1\}^{t\times t}$ that serves as a mask to determine whether there will be new residual connections [68] between pairs of encoders and decoders. More concretely, $\mathbf{A}_{ij}=1$ indicates that there exists a skip connection from $e_{i}$ to $d_{j}$ and $\mathbf{A}_{ij}=0$ means that there is not such a skip connection. As the shapes of feature maps across different network levels can vary significantly (e.g., $64\times 64$ for the output of $e_{i}$ and $256\times 256$ for the input of $d_{j}$), we introduce connection scale factors to tackle this inconsistency and decompose all the possible scale factors into a series of $2\times$ upsampling operations or downsampling operations with shared weights. By allocating 0 or $1$ to each of the $t^{2}$ bits in A, we broaden the search space of skip connection patterns as there are $2^{t^{2}}$ possibilities altogether and we only need to sample a portion of them.
@@ IV-B2 Optimal Model Selection
We build up our model search space $\mathcal{M}$ by combining the possibilities of the aforementioned upsampling modules and skip connection patterns. We assume that the size of our model search space is $n$ and the candidates inside it are denoted as $\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$. We first sample the latent code $\mathbf{z_{0}}$ from the Gaussian distribution and freeze its values on all the models for fair comparison. We then traverse the model search space and calculate the initial gradient matching loss of each individual $G_{r}\,(1\leq r\leq n)$:
@@ IV-C Stage 2: Private Batch Recovery via the Optimal Model
[…]
3:Build up the model search space: $\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$ with the parameters $\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$
[…]
@@ V-A Experimental Setup
Evaluation Settings. We test our method on CIFAR-10 [69] and ImageNet [22] with the resolutions of $32\times 32$ and $256\times 256$. Unlike many previous methods [18, 12] that scale down the ImageNet images to $64\times 64$, here we emphasize that we adopt the high-resolution version of ImageNet. Thus, our settings are more rigorous and realistic. Following previous gradient inversion works [9, 13], we adopt ResNet-$18$ [68] as our global model and utilize the same preprocessing procedures. We build up the search space and randomly generate the alternative models by arbitrarily changing the options of upsampling modules and skip connection patterns.
@@ V-D Further Analysis
TABLE IX: Attack performance when increasing the training iterations for the search metric on ImageNet ($256\times 256$) with the default batch size $B=4$. Note that the training iteration of 0 indicates the use of our training-free search metric.
@@
Ablation Study on Search Metric Training Iteration. GI-NAS utilizes the initial gradient matching loss as the training-free search metric, which implies that its training iteration for each candidate is 0. To explore the rationality behind this design, we gradually increase the training iterations for the search metric. From Table IX, we conclude that further increasing the training iterations cannot improve the attacks and will instead increase the time costs. Hence, utilizing the initial gradient matching loss to find the optimal models is reasonable and computationally efficient.
@@
Dataset Model Random Seed Initial Gradient Matching Loss ($\times 10^{-3}$) $\sigma_{loss}$ ($\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{C1}$ $Seed_{C1}$ 8.5647 0.0012 28.7283 0.1211 $Seed_{C2}$ 8.5648 28.8082 $Seed_{C3}$ 8.5645 28.5703 $G_{C2}$ $Seed_{C1}$ 8.4321 0.0006 33.5848 0.2581 $Seed_{C2}$ 8.4323 33.4631 $Seed_{C3}$ 8.4322 33.9584 $G_{C3}$ $Seed_{C1}$ 7.9501 0.0015 49.4290 0.2936 $Seed_{C2}$ 7.9499 49.1171 $Seed_{C3}$ 7.9502 49.7039 ImageNet $G_{I1}$ $Seed_{I1}$ 21.0167 5.6866 22.9013 0.0329 $Seed_{I2}$ 22.1401 22.8912 $Seed_{I3}$ 21.4250 22.8398 $G_{I2}$ $Seed_{I1}$ 24.4600 2.2089 21.1475 0.3040 $Seed_{I2}$ 24.0722 21.6449 $Seed_{I3}$ 24.0828 21.6989 $G_{I3}$ $Seed_{I1}$ 18.5319 6.0290 23.5278 0.1884 $Seed_{I2}$ 18.3819 23.2974 $Seed_{I3}$ 17.4208 23.6709
@@
Dataset Model Random Seed Initial Gradient Matching Loss ($\times 10^{-3}$) $\sigma_{loss}$ ($\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{C4}$ $Seed_{C4}$ 14.5338 0.0018 26.4506 0.1925 $Seed_{C5}$ 14.5335 26.2082 $Seed_{C6}$ 14.5336 26.0703 $G_{C5}$ $Seed_{C4}$ 13.4751 0.0018 30.1240 0.2675 $Seed_{C5}$ 13.4753 30.2045 $Seed_{C6}$ 13.4755 30.6222 $G_{C6}$ $Seed_{C4}$ 13.2704 0.0015 31.3139 0.0566 $Seed_{C5}$ 13.2701 31.4256 $Seed_{C6}$ 13.2702 31.3540 ImageNet $G_{I4}$ $Seed_{I4}$ 23.1931 4.6280 21.8376 0.1164 $Seed_{I5}$ 23.6335 22.0556 $Seed_{I6}$ 22.7083 21.8760 $G_{I5}$ $Seed_{I4}$ 19.5264 1.1102 23.6033 0.1889 $Seed_{I5}$ 19.5677 23.2439 $Seed_{I6}$ 19.7360 23.3226 $G_{I6}$ $Seed_{I4}$ 25.9479 3.6054 20.4730 0.1655 $Seed_{I5}$ 25.3497 20.8007 $Seed_{I6}$ 25.9975 20.6776
@@
Preferences of NAS Outcomes. To figure out the implications behind NAS preferences, we trace back the searched optimal models from Section V-C and analyze their averaged architectural statistics in Table XIII. Note that the bicubic operation takes more neighbourhood pixels than the bilinear operation ($16>4$) and is thus considered to be more complex. Besides, the pixel shuffle operation rearranges lots of pixels and is also considered to be more complicated than the bilinear operation. From Table XIII, we observe that the optimal models are largely dependent on recovery difficulties. Generally, batches with larger sizes or under defenses are more difficult to recover. Compared to the default settings (highlighted in gray color) that are with the batch size $B=4$ and under no defenses, the more difficult settings would prefer models with higher complexity (e.g., more skip connections, larger kernel sizes, more complex operations). The complexity gains $\Delta_{connection}$ and $\Delta_{upsampling}$ are often greater than 0 when the settings are $B>4$ or under defenses. Easier batches prefer simpler, lighter architectures and harder batches prefer more complex, heavier architectures. This might be used to further improve the efficiency. By considering the difficulties ahead of time, we can have a better estimation of the complexity of optimal models. Thus, we only need to focus on models with a certain type of complexity and skip other unrelated models even when there is not enough time for NAS.